OpenStudy (nannie):

PLEASE HELP!!! Solve the equation w=Cr^2 for r. (the weight of an object follows this equation: , where C is a constant, and r is the distance that the object is from the center of Earth.)

OpenStudy (anonymous):

r=(w/C)^0.5

OpenStudy (anonymous):

\[r = \sqrt{\frac{w}{C}}\]